OpenStudy (anonymous):

1000 x 10^-10 helppp? walkthrough?

OpenStudy (mathstudent55):

What is the question? You have a number written in scientific notation. What are you trying to do?

OpenStudy (anonymous):

noo its not scientific notation. its actually like seperate numbers being multiplied. "An x-ray length is 10^-10 m. An ultraviolet wave is ten times the wavelength of an x-ray. Find the wavelength of ultraviolet waves."

OpenStudy (mathstudent55):

If it's ten times, why are you multiplying by 1000?

OpenStudy (anonymous):

\[10^-10\] is 10 * -10 * -10 * -10 * -10 * -10 * -10 * -10 * -10 * -10 * -10.

OpenStudy (anonymous):

Oh wait, I think I messed up.

OpenStudy (mathstudent55):

10 * 10^-10 = 10 / 10^10 = 10^(1-10) = 10^-9

OpenStudy (mathstudent55):

Remember division of powers with the same base: a^m/a^n = a^(m - n)

OpenStudy (anonymous):

It's a thousand times 10 to the 10th power, though, math. It wouldn't be 10 * 10^-10, would it?

OpenStudy (mathstudent55):

I still don't understand where the 1000 times comes from. The problem states a length of a certain wave is 10^-10 m, and the other length is 10 times larger. Where is the 1000 coming from?

OpenStudy (anonymous):

i meant 1000 times larger

OpenStudy (mathstudent55):

If you do want to do 1000 * 10^-10, then rewrite the 1000 as a power of 10: 1000 * 10^-10 = 10^3 * 10^-10 = 10^-7 The rule of multiplying powers with the same base is: a^m * a^n = a*(m + n)

OpenStudy (mathstudent55):

Oh, I see, it's 1000 time larger. Got it.

OpenStudy (anonymous):

yeah so what do i do